# Reversal Blessing: Thinking Backward May Outpace Thinking Forward in Multi-choice Questions

### v2

@@ 4 Controlled Simulation Study
The three hypotheses discussed in Section 3 are intricately entwined in actual MCQs, making it challenging to disentangle them. To better investigate the hypotheses explaining the optimal direction for MCQs, we conducted a meticulously controlled simulation study (Figure 4) focused on 4-digit multiplication. The L2R and R2L models were initialized from scratch and exclusively trained on this simulation dataset to eliminate any potential confounding factors. All data instances share the same format and length, removing the calibration effect from the analysis and allowing us to concentrate on computability and conditional entropy.
@@
We conduct two types of simulation experiments: Forward Multiplication (Forward X) and Reverse Multiplication (Reverse X). In Forward X, each training instance was represented as $m\times n=p$, where $m,n\in\{0,\ldots,10^{4}\}$ and $p\in\{0,\ldots,10^{8}\}$. The formatting included spaces between digits and mathematical operators to ensure a consistent single tokenization for both L2R and R2L models. In Reverse X, the multiplication was in reverse order, such as $p=m\times n$. For each simulation type, L2R and R2L models were trained with a 2B model size with 1 epoch on all $10^{8}$ non-repeating equations except 1,000 test examples, totaling to almost 3.2B tokens.
The model performance was assessed using the held-out test set with 1,000 examples. These examples were converted into a multiple-choice format consisting of 4 choices (Figure 4). Other than the correct answer, the remaining three hard-negative options were created by altering a single digit in the correct answer to a random other digit, at a random position. The presenting order of the four choices are then randomly shuffled. We augmented the test set 10 times and calculated the average accuracy and conditional entropy.
@@ Results
The results are presented in Table 2. In Forward X scenarios, L2R models demonstrate significantly higher accuracies than R2L(m,n) models, with correspondingly lower conditional entropy and training loss. This observation aligns with our hypothesis in Section 3. Conversely, in Reverse X scenarios, the R2L model outperforms the L2R(m,n) model. The training and test performance gaps are minimal.
@@
On the other hand, comparing L2R with R2L(m), where theoretical conditional entropy equal 0, L2R maintains superiority, indicating that computability likely remains as a key factor. For the Reverse X task, the accuracy gap between R2L and L2R(n) is smaller than the accuracy gap between L2R(m,n) and L2R(n), suggesting that the conditional entropy may explain more of the performance gap than the computability.
Notably, models achieve higher accuracies on Reverse X compared to their Forward X counterparts, despite similar training loss and conditional entropy values. This disparity could probably be attributed to the closer proximity of choices in Forward X, which inherently increases task difficulty. We provide additional discussion and analysis comparing Forward X and Reverse X in Appendix E.
@@ 5 Related Work
Berglund et al. (2023) first investigates the "reversal curse" in LLMs, which refers to the phenomenon where models trained on forward text data struggle to perform well on inverse search tasks. Allen-Zhu and Li (2023a) further discusses this issue and proposes that augmentation during the pretraining stage can help bridge the knowledge extraction performance gap in reverse entity mapping. In a similar vein, Golovneva et al. (2024) suggest training a unified model that combines text data with augmented reversed or partially reversed data can mitigate the reversal curse. These studies imply that autoregressively-trained language models tend to have a linear and unidirectional thinking process, and certain types of augmentation can faciliate the model in making complex connections between pieces of learned information to enable more intricate cross-referencing. Our research also demonstrates that the autoregressive nature of LLMs may introduce inductive biases rooted from the pretraining corpus. Instead of focusing on the "reversal curse," we suggest that knowledge extraction and reasoning may be more straightforward in the direction with lower conditional entropy.
Order of Reasoning
@@
Previous works have also been exploring the reasoning order’s impact to the reasoning performance. Vinyals et al. (2015) first demonstrates that the sequence in which input and output data are organized significantly impacts the performance of sequence-to-sequence models and propose to search over possible orders during training to manage unstructured output sets. Recently, Papadopoulos et al. (2024) reveals a surprisingly consistent lower log-perplexity when predicting in L2R versus R2L, despite theoretical expectations of symmetry. The authors attributes this asymmetry to factors like sparsity and computational complexity. We also observe this difference yet we have another hypothesis rationale beyond theirs. Zhang-Li et al. (2024) shows that by reversing the digit order, prioritizing the least significant digit can improve LLMs’s performance on arithmetic, which aligns with our findings in section 4.
@@
Previous studies on sequence modeling have also delved into relaxing the conventional “left-to-right” autoregressive dependencies, primarily to facilitate rapid parallel generation (Gu et al., 2018; Ghazvininejad et al., 2019; Gu and Kong, 2021; Zhang et al., 2020) and non-monotonic sequential generation (Welleck et al., 2019; Gu et al., 2019). Text diffusion has recently emerged as a promising approach in terms of planning and controllability (Li et al., 2022; Zhang et al., 2023; Gong et al., 2024). It has shown to be more effective than LLM than language model (LLM), particularly for tasks that require bidirectional reasoning strategies such as sudoku and countdown games (Ye et al., 2024).
Multiple-Choice Questions (MCQs) for LLM evaluation
@@
MCQs have been widely used for evaluation LLM’s reasoning and knowledge extraction abilities. Zheng et al. (2023) demonstrates that LLMs exhibit a selection bias in MCQs, favoring certain option positions, and introduces a debiasing method to mitigate this issue. Pezeshkpour and Hruschka (2023) examines how LLMs’ performance on MCQs is influenced by the order of answer options, finding that reordering can lead to huge performance variations. Ghosal et al. (2022) proposes reframing MCQs as a series of binary classifications, demonstrating that this approach significantly improves performance across various models and datasets. Li et al. (2024b) highlights issues like positional biases and discrepancies compared to long-form generated responses, when using MCQs in evaluating LLMs. Wiegreffe et al. (2024) discovers that the prediction of specific answer symbols is primarily attributed to a single middle layer’s multi-head self-attention mechanism, with subsequent layers increasing the probability of the chosen answer in the model’s vocabulary space. In contrast to the previous work, our work first shows the connection between the preferred reasoning direction and the direction that has lower conditional entropy in MCQ evaluations.
@@ 6 Conclusion
In this work, we investigated the potential benefits of R2L factorization in language modeling, focusing on MCQs. Through extensive experimentation with models of varying sizes and training datasets, we demonstrated that R2L factorization can outperform traditional L2R approaches in specific MCQs. Our analysis revealed that the effectiveness of each factorization direction may be intrinsically linked to several factors including calibration, computability, and conditional entropy of the downstream task distribution, with lower conditional entropy yielding better performance. We disentangle and validate these factors through controlled simulation studies using arithmetic tasks. These findings may suggest the potential for future language model development, by revealing the knowledge extraction and reasoning machinery of LLM and suggesting that alternative factorizations deserve serious consideration in model design. Future work could explore additional factorization strategies beyond L2R and R2L, investigate applications to other types of language tasks, and develop more sophisticated methods for combining different factorizations.
@@ Acknowledgment
Limitation
Our work has several limitations. While MCQs provide a controlled evaluation setting, they represent only a subset of language understanding tasks, and the applicability of our findings to other formats like open-ended generation remains unexplored. We are currently extending our investigation to include generative tasks and dialogue systems to validate the broader applicability of our directional factorization hypothesis. Our theoretical framework around conditional entropy, while supported by empirical observations, lacks formal proofs and relies on estimated conditional entropy as an imperfect proxy. The simulation studies focused primarily on arithmetic operations with well-defined properties, which may not fully generalize to more complex language understanding and reasoning scenarios. New controlled experiments yet to be designed with increasingly complex reasoning chains and various answer length that better mirror real-world language understanding tasks. Our experiments were limited to models in the 2B-8B parameter range, and the relationships we observed might vary with different model scales or architectures.
@@ Appendix A Model Architecture and Training Hyperparameters
Parameter 2B Model 8B Model
Vocabulary Size 128,256 128,256
Number of Layers 24 32
Number of Heads 16 32
Number of KV Heads 8 8
Embedding Dimension 2,048 4,096
Max Sequence Length 2,048 8,192
Intermediate Dimension 7,168 14,336
RoPE Base 500,000 500,000
Appendix B Comparing Three Paradigms for Reverse Thinking
@@
Appendix C Experimental Details on Computing the Conditional Entropy in MCQs
@@
Appendix D Analysis of How Reverse Thinking Reduces Surface Form Competition in MCQ
@@
Appendix E Additional Results for Simulation Study

### v3

@@ 4 Controlled Simulation Study
The three hypotheses discussed in Section 3 are intricately entwined in actual MCQs, making it challenging to disentangle them. To better investigate the hypotheses explaining the optimal direction for MCQs, we conducted a meticulously controlled simulation study (Figure 4) focused on 4-digit multiplication. Although the arithmetic dataset is different from the real language modeling datasets, this simulation study can be a good controlled experiment to understand the phenomenon we observed in Section 3, as we are investigating the underlying principle of the model with different factorizations regardless of the dataset. The L2R and R2L models were initialized from scratch and exclusively trained on this simulation dataset to eliminate any potential confounding factors. All data instances share the same format and length, removing the calibration effect from the analysis and allowing us to concentrate on computability and conditional entropy.
@@
We conduct two types of simulation experiments: Forward Multiplication (Forward X) and Reverse Multiplication (Reverse X). In Forward X, each training instance was represented as $m\times n=p$, where $m,n\in\{0,\ldots,10^{4}\}$ and $p\in\{0,\ldots,10^{8}\}$. The formatting included spaces between digits and mathematical operators to ensure a consistent single tokenization for both L2R and R2L models. In Reverse X, the multiplication was in reverse order, such as $p=m\times n$. For each simulation type, L2R and R2L models were trained with a 2B model size with 1 epoch on all $10^{8}$ non-repeating equations except 1,000 test examples, totaling almost 3.2B tokens.
The model performance was assessed using the held-out test set with 1,000 examples. These examples were converted into a multiple-choice format consisting of 4 choices (Figure 4). Other than the correct answer, the remaining three hard negative options were created by altering a single digit in the correct answer to a random other digit, at a random position. The presenting order of the four choices are then randomly shuffled. We augmented the test set 10 times and calculated the average accuracy and conditional entropy.
@@ Results
The results are presented in Table 2. In Forward X scenarios, L2R models demonstrate higher accuracies than R2L(m,n) models, with correspondingly lower conditional entropy and training loss. This observation aligns with our hypothesis in Section 3. Conversely, in Reverse X scenarios, the R2L model outperforms the L2R(m,n) model. The training and test performance gaps are minimal.
@@
On the other hand, comparing L2R with R2L(m), where theoretical conditional entropy equals 0, L2R maintains superiority, indicating that computability likely remains as a key factor. For the Reverse X task, the accuracy gap between R2L and L2R(n) is smaller than the accuracy gap between L2R(m,n) and L2R(n), suggesting that the conditional entropy may explain more of the performance gap than the computability.
Notably, models achieve higher accuracies on Reverse X compared to their Forward X counterparts, despite similar training loss and conditional entropy values. This disparity could probably be attributed to the closer proximity of choices in Forward X, which inherently increases task difficulty. We provide additional discussion and analysis comparing Forward X and Reverse X in Appendix F.
@@ 5 Related Work
Reversal Curse Berglund et al. (2023) first investigates the "reversal curse" in LLMs, which refers to the phenomenon where models trained on forward text data struggle to perform well on inverse search tasks. Allen-Zhu & Li (2023a) further discusses this issue and proposes that augmentation during the pretraining stage can help bridge the knowledge extraction performance gap in reverse entity mapping. In a similar vein, Golovneva et al. (2024) suggest training a unified model that combines text data with augmented reversed or partially reversed data can mitigate the reversal curse. These studies imply that autoregressively-trained language models tend to have a linear and unidirectional thinking process, and certain types of augmentation can facilitate the model in making complex connections between pieces of learned information to enable more intricate cross-referencing. Our research also demonstrates that the autoregressive nature of LLMs may introduce inductive biases rooted from the pretraining corpus. Instead of focusing on the "reversal curse," we suggest that knowledge extraction and reasoning may be more straightforward in the direction with lower conditional entropy.
@@
Order of Reasoning Previous work has also explored the reasoning order’s impact on reasoning performance. Vinyals et al. (2015) first demonstrates that the sequence in which input and output data are organized significantly impacts the performance of sequence-to-sequence models and propose to search over possible orders during training to manage unstructured output sets. Recently, Papadopoulos et al. (2024) reveals a surprisingly consistent lower log-perplexity when predicting in L2R versus R2L, despite theoretical expectations of symmetry. The authors attribute this asymmetry to factors like sparsity and computational complexity. We also observe this difference yet we have another hypothesis rationale beyond theirs. Zhang-Li et al. (2024) shows that by reversing the digit order, prioritizing the least significant digit can improve LLMs’s performance on arithmetic, which aligns with our findings in section 4.
@@
Previous studies on sequence modeling have also delved into relaxing the conventional “left-to-right” autoregressive dependencies, primarily to facilitate parallel generation (Gu et al., 2018; Ghazvininejad et al., 2019; Gu & Kong, 2021; Zhang et al., 2020) and non-monotonic generation (Welleck et al., 2019; Gu et al., 2019). Text diffusion has recently emerged as a promising approach in terms of planning and controllability (Li et al., 2022; Zhang et al., 2023; Gong et al., 2024). It has shown to be more effective than language models (LLMs), particularly for tasks that require bidirectional reasoning strategies such as sudoku and countdown games (Ye et al., 2024). Alternatively, the Belief State Transformer (BST) (Hu et al., 2025) enhances sequence modeling by using both prefix and suffix inputs to predict subsequent and preceding tokens, effectively capturing a compact belief state for improved goal-conditioned decoding and test-time inference. In contrast, our work primarily investigates the optimal reasoning order for non-generative tasks requiring structured inference.
@@
Multiple-Choice Questions (MCQs) for LLM evaluation MCQs have been widely used for evaluating LLM’s reasoning and knowledge extraction abilities. Zheng et al. (2023) demonstrates that LLMs exhibit a selection bias in MCQs, favoring certain option positions, and introduces a debiasing method to mitigate this issue. Pezeshkpour & Hruschka (2023) examines how LLMs’ performance on MCQs is influenced by the order of answer options, finding that reordering can lead to huge performance variations. Ghosal et al. (2022) proposes reframing MCQs as a series of binary classifications, demonstrating that this approach significantly improves performance across various models and datasets. Li et al. (2024b) highlights issues like positional biases and discrepancies compared to long-form generated responses, when using MCQs in evaluating LLMs. Wiegreffe et al. (2024) discovers that the prediction of specific answer symbols is primarily attributed to a single middle layer’s multi-head self-attention mechanism, with subsequent layers increasing the probability of the chosen answer in the model’s vocabulary space. In contrast to the previous work, our work first shows the connection between the preferred reasoning direction and the direction that has lower conditional entropy in MCQ evaluations.
6 Limitations
Task Scope and Generalizability While our findings demonstrate the benefits of R2L factorization in some MCQs, the applicability to other NLP tasks remains unexplored. Our focus on MCQs was intentional to provide a controlled setting for studying factorization effects, as evaluation on generative tasks is not as straightforward. While exploring similar observations in other generative tasks is a valuable avenue for future research, in this paper we limit our discussion to MCQs as a lens to investigate the machinery of LLMs. Additionally, our experiments were limited to models in the 2B-8B parameter range. Due to resource constraints, we were not able to pretrain a SOTA-scale LLM from scratch; future work should explore whether our findings generalize to SOTA models.
Theoretical Framework We propose three hypotheses (calibration, computability, and conditional entropy) to explain different factorization directions’ effectiveness, but our analysis in natural language datasets remains qualitative as it is challenging to measure these factors in real language. We used synthetic arithmetic tasks to better understand factors contributing to the L2R and R2L performance gap. Future work should develop rigorous quantitative measures and establish formal theoretical connections between these factors through more sophisticated controlled experiments.
Experimental and Model Design Our evaluation methodology, particularly choosing between $p(q|a_{i})$ and $p(a_{i}|q)$ paradigms, may introduce confounding factors that make isolating pure factorization effects challenging. Future work should develop more robust frameworks that better disentangle these effects from task reformulation. While we focus on standalone L2R and R2L models, incorporating both factorizations through a unified model may further improve performance.
@@ 7 Conclusion
In this work, we investigated what makes the preferred thinking direction for LLMs in multiple-choice questions. Through extensive experimentation with models of varying sizes and training datasets, we discovered the surprising finding that R2L factorization can outperform traditional L2R approaches in specific MCQ tasks (4 out of 11 evaluated benchmarks). Our analysis revealed that the effectiveness of each factorization direction may be intrinsically linked to several factors including calibration, computability, and conditional entropy of the downstream task distribution, with lower conditional entropy yielding better performance. We disentangled and validated these factors through controlled simulation studies using arithmetic tasks.
The core contribution of this work lies in bringing this phenomenon to the community’s attention and initiating analysis into the underlying factors that might explain why a particular thinking direction is preferred in specific task domains. These findings may suggest the potential for future language model development by revealing the knowledge extraction and reasoning machinery of LLMs and suggesting that alternative factorizations deserve serious consideration in model design. Future work could explore additional factorization strategies beyond L2R and R2L, investigate applications to other types of language tasks, and develop more sophisticated methods for combining different factorizations.
@@ Appendix A Model Architecture and Training Hyperparameters
Parameter 1.5B Model 2B Model 4B Model 8B Model
Vocabulary Size 128,256 128,256 128,256 128,256
Number of Layers 18 24 24 32
Number of Heads 16 16 16 32
Number of KV Heads 16 8 8 8
Embedding Dimension 2048 2,048 3072 4,096
Max Sequence Length 2,048 2048 2,048 8,192
Intermediate Dimension 5,632 7,168 11,264 14,336
RoPE Base 500,000 500,000 500,000 500,000
Appendix B Results with Different Model Sizes
We conducted additional experiments to explore performance trends across a broader range of model scales, training 1B and 4B parameter models (R2L and L2R) to complement our existing 2B and 8B models. The full results, presented in Table 4, consistently reinforce our main claim.
Table 4: Results with different model sizes.
EDU-1.5B EDU-2B EDU-4B EDU-8B
L2R R2L L2R R2L L2R R2L L2R R2L
LogiQA 28.10 32.65 27.96 31.49 26.58 32.72 29.95 31.03
OpenbookqQA 38.00 41.20 42.40 44.40 43.00 44.60 45.00 48.40
TruthfulQA 23.01 28.86 24.36 28.76 22.40 29.00 24.97 31.70
CommensenseQA 41.44 44.22 42.92 45.13 50.70 43.40 39.15 44.96
Social_IQA 41.35 41.65 42.78 42.22 42.99 40.00 44.58 43.50
AI2_arc 58.63 51.8 60.65 52.31 63.14 53.95 68.29 56.22
HellaSwag 56.46 43.38 60.57 44.34 65.93 46.38 71.60 49.22
MathQA 25.52 24.05 26.80 24.86 27.33 24.08 28.77 25.33
MMLU 33.30 33.40 34.57 34.35 36.70 35.16 38.90 37.11
PIQA 73.23 59.19 74.48 57.13 76.00 58.76 77.80 59.14
Winogrande 58.56 51.77 60.93 54.85 61.80 54.50 65.75 54.70
Specifically, we observe a clear scaling trend: for the tasks where reverse thinking excels (LogiQA, OpenbookQA, TruthfulQA, and Commonsense QA), the performance advantage of R2L over L2R grows as the model size increases. Although we did not scale beyond 8B parameters due to engineering constraints, these extended results demonstrate the robustness of our findings. This trend strongly suggests that the impact of alternative factorizations is unlikely to diminish with even larger models; on the contrary, their benefits may become more significant.
Appendix C Comparing Three Paradigms for Reverse Thinking
@@
Appendix D Experimental Details on Computing the Conditional Entropy in MCQs
@@
Appendix E Analysis of How Reverse Thinking Reduces Surface Form Competition in MCQ
@@
Appendix F Additional Results for Simulation Study
@@
Appendix G Statistical Testing Results
We presents detailed statistical comparisons between L2R and R2L reasoning approaches across four tasks where R2L outperforms L2R: LogiQA, OpenBookQA, TruthfulQA, and CommonsenseQA. Metrics include mean accuracy, standard deviation, and the results of a paired t-test comparing the two approaches. All results were computed using bootstrap sampling with five replicates, each resampled at 80% with replacement, providing a robust estimation of performance variance. Notably, R2L consistently outperformed L2R, with statistically significant improvements ($p$-value < 0.05) observed across all tasks. The largest relative performance gain appeared in LogiQA, where R2L achieved a mean accuracy increase from 27.04% (L2R) to 32.56% (R2L) with an extremely significant $p$-value. OpenBookQA also exhibited a clear improvement, rising from 42.80% to 44.40%. TruthfulQA and CommonsenseQA saw meaningful gains as well, with $p$-values of 0.0072 and 0.0471, respectively, indicating reliable improvements beyond random fluctuation. Overall, these results underscore the effectiveness of reverse thinking (R2L) for improving reasoning tasks.
Table 6: Paired t-test of bootstrap samples (5 replicates, 80% resampled with replacement).
Task L2R Mean L2R Std R2L Mean R2L Std $p$-value
LogiQA 27.04 0.00 32.56 0.0045 0.0000
OpenBookQA 42.80 0.00 44.40 0.0000 0.0000
TruthfulQA 25.08 1.16 28.93 0.80 0.0072
CommonsenseQA 42.17 1.34 44.68 1.70 0.0471